# Bayesian optimization of electron energy from laser wakefield accelerator

### v2

@@ VI.2 Guiding in preformed plasma channel with matched radius
Panels (a) and (b) of Fig. 5 show the spatial and spectral characteristics, respectively, of the output electron beam obtained from $B_{17}$ – the PIC simulation with parameters closest to the optimum found by BO in a preformed plasma channel (parameters are listed in Tab. 2). The total charge of electrons with energy above $1\,\mathrm{MeV}$ is now $\approx 149\,\mathrm{pC}$, nearly four times higher compared to the optimal case in a uniform-density plasma. The angular distribution of the electron beam shows a FWHM divergence of $\approx 18\,\mathrm{mrad}$. Again, no effort was made to optimize these characteristics.
For the optimal parameters, the maximum electron energy according to Eq. (24) is $\mathcal{E}_{e,\mathrm{max}}\approx 70\,\mathrm{MeV}$, and the corresponding acceleration distance according to Eq. (25) is $l_{\mathrm{acc}}\approx 113\,\upmu\mathrm{m}$. The discrepancy between the theoretical prediction and simulation results can be attributed to the same factors as in the case of self-guiding in a uniform-density plasma. Furthermore, the increase in electron energy and acceleration distance may be related to the mechanism described in Ref. 56: at later stages of the interaction, there is a significant overlap between the accelerated electrons and the red-shifted field of the laser pulse [see Fig. 5(a)], which may lead to the onset of direct laser acceleration [57]. This additional mechanism could enhance the final energy gain of the accelerated electrons, complementing the primary process of LWFA.
@@ VII Discussion on scaling
As shown in Sec. II, LWFA operated in the matched regime can be described using only two parameters, $n_{e}$ and $\tau_{0}$; alternatively, these parameters can be expressed in dimensionless form as $\mathcal{P}^{*}\coloneq\mathcal{P}_{0}/\mathcal{P}_{\mathrm{cr}}$ and $\tau^{*}\coloneq\tau_{0}\omega_{p}$. In general, $\mathcal{P}^{*}$ and $\tau^{*}$ may be functions that depend weakly on the laser energy. Treating them as constant for now, and disregarding the effects mentioned in the previous paragraph, we may derive the scaling of optimal LWFA input parameters, i.e., plasma density, pulse duration (or power), amplitude, and spot size, all expressed in terms of laser energy:
@@
$\approx 0.49\,\left(\tau^{*}\right)^{2/3}\left(\mathcal{P}^{*}\right)^{2/3}\left(\frac{\mathcal{E}_{0}}{\overline{\mathcal{E}}}\right)^{-2/3},$
@@
$\approx 0.23\,\left(\tau^{*}\right)^{2/3}\left(\mathcal{P}^{*}\right)^{-1/3}\left(\frac{\mathcal{E}_{0}}{\overline{\mathcal{E}}}\right)^{1/3},$
@@
$\approx 2.06\,\left(\tau^{*}\right)^{-2/3}\left(\mathcal{P}^{*}\right)^{1/3}\left(\frac{\mathcal{E}_{0}}{\overline{\mathcal{E}}}\right)^{2/3},$
@@
$\approx 0.65\,\left(\tau^{*}\right)^{-1/3}\left(\mathcal{P}^{*}\right)^{-1/6}\left(\frac{\mathcal{E}_{0}}{\overline{\mathcal{E}}}\right)^{1/3}.$
@@
$\frac{\mathcal{E}_{e,\mathrm{max}}}{m_{e}c^{2}}$ $\approx 2.08\left(\frac{2\sqrt{\pi\ln{2}}}{\tau^{*}}\right)^{2/3}\left(\frac{\mathcal{E}_{0}}{\overline{\mathcal{E}}}\right)^{2/3}$ (33)
$\approx 4.28\left(\tau^{*}\right)^{-2/3}\left(\frac{\mathcal{E}_{0}}{\overline{\mathcal{E}}}\right)^{2/3},$
@@
$\approx 0.08\frac{\lambda_{0}}{r_{e}}\left(\tau^{*}\right)^{-1/3}\left(\mathcal{P}^{*}\right)^{1/6}\left(\frac{\mathcal{E}_{0}}{\overline{\mathcal{E}}}\right)^{1/3}.$
In the case of laser self-guiding in a uniform-density plasma (Sec. VI.1), the optimal case found in this work corresponds to $\mathcal{P}^{*}\approx 2.31$ and $\tau^{*}\approx 2.28$. Using Eqs. (28)–(32) to rescale the optimal parameters to, e.g., a $1\,\mathrm{kJ}$ laser pulse, we obtain $n_{e}\approx 1.56\times 10^{16}\,\mathrm{cm^{-3}}$, $\tau_{0}\approx 327\,\mathrm{fs}$, $w_{0}\approx 140\,\upmu\mathrm{m}$, and $a_{0}\approx 2.65$. For such a hypothetical LWFA setup, Eqs. (33) and (34) predict a maximum electron energy of $\mathcal{E}_{e,\mathrm{max}}\approx 134\,\mathrm{GeV}$ and a corresponding acceleration length of $l_{\mathrm{acc}}\approx 9.8\,\mathrm{m}$.
In the case of laser guiding in a preformed plasma channel with matched radius (Sec. VI.2), the optimal case corresponds to $\mathcal{P}^{*}\approx 2.06$ and $\tau^{*}\approx 1.88$. Considering again a $1\,\mathrm{kJ}$ laser pulse, the optimal parameters translate to $n_{e}\approx 1.27\times 10^{16}\,\mathrm{cm^{-3}}$, $\tau_{0}\approx 299\,\mathrm{fs}$, $w_{0}\approx 152\,\upmu\mathrm{m}$, $a_{0}\approx 2.54$, and $R_{\mathrm{ch}}\approx 140\,\upmu\mathrm{m}$. Such a setup would yield an electron beam with a cut-off energy of $\mathcal{E}_{e,\mathrm{max}}\approx 161\,\mathrm{GeV}$ over an acceleration distance of $l_{\mathrm{acc}}\approx 11\,\mathrm{m}$.
Panels (a)–(f) of Fig. 6 show the scaling of optimal plasma density, laser pulse duration, laser power, beam waist, electron beam cut-off energy, and acceleration length, respectively, all expressed as functions of laser energy $\mathcal{E}_{0}$, according to Eqs. (28), (29), (30), (32), (33), and (34). Results are shown for both cases: laser self-guiding in a uniform-density plasma and guiding in a preformed plasma channel with matched radius.
@@ VIII Conclusion
To interpret the simulation results quantitatively, we derive novel analytical expressions for predicting the maximum electron energy and the corresponding acceleration length, accounting for the effects of laser pulse diffraction and energy depletion. The model shows good agreement with the simulation results.
@@ Appendix A Simulation parameters
##\## $n_{e}\left(10^{19}\mathrm{cm^{-3}}\right)$ $\tau_{0}\left(\mathrm{fs}\right)$ $w_{0}\left(\mathrm{\upmu m}\right)$ $a_{0}$ $\mathcal{E}_{e,\mathrm{max}}\left(\mathrm{MeV}\right)$ $l_{\mathrm{acc}}\left(\mathrm{\upmu m}\right)$
@@
Table 1: Parameters of PIC simulations with a uniform-density plasma. Electron density, $n_{e}$, pulse duration, $\tau_{0}$, beam waist, $w_{0}$, and strength parameter, $a_{0}$, prescribed in PIC simulations $A_{1}$ $A_{18}$ as well as the corresponding cut-off electron energy $\mathcal{E}_{e,\mathrm{max}}$ acquired over acceleration distance $l_{\mathrm{acc}}$.
##\## $n_{e}\left(10^{19}\mathrm{cm^{-3}}\right)$ $\tau_{0}\left(\mathrm{fs}\right)$ $w_{0}\left(\mathrm{\upmu m}\right)$ $a_{0}$ $R_{\mathrm{ch}}\left(\mathrm{\upmu m}\right)$ $\mathcal{E}_{e,\mathrm{max}}\left(\mathrm{MeV}\right)$ $l_{\mathrm{acc}}\left(\mathrm{\upmu m}\right)$
@@
Table 2: Parameters of PIC simulations with a preformed plasma channel of matched radius. Electron density, $n_{e}$, pulse duration, $\tau_{0}$, beam waist, $w_{0}$, strength parameter, $a_{0}$, and the channel radius, $R_{\mathrm{ch}}$, prescribed in PIC simulations $B_{1}$ $B_{18}$ as well as the corresponding cut-off electron energy $\mathcal{E}_{e,\mathrm{max}}$ acquired over acceleration distance $l_{\mathrm{acc}}$.

### v3

@@ VI.2 Guiding in preformed plasma channel with matched radius
Panels (a) and (b) of Fig. 5 show the spatial and spectral characteristics, respectively, of the output electron beam obtained from $B_{17}$ – the PIC simulation with parameters closest to the optimum found by BO in a preformed plasma channel (parameters are listed in Tab. 2). The total charge of electrons with energy above $1\,\mathrm{MeV}$ is now $\approx 149\,\mathrm{pC}$, nearly four times higher compared to the optimal case in a uniform-density plasma. The angular distribution of the electron beam shows a FWHM divergence of $\approx 18\,\mathrm{mrad}$ and the efficiency of energy transfer from the laser to the accelerated electron beam is $\approx 31.9\,\%$. Again, no effort was made to optimize these characteristics.
For the optimal parameters, the maximum electron energy according to Eq. (25) is $\mathcal{E}_{e,\mathrm{max}}\approx 70\,\mathrm{MeV}$, and the corresponding acceleration distance according to Eq. (26) is $l_{\mathrm{acc}}\approx 113\,\upmu\mathrm{m}$. The discrepancy between the theoretical prediction and simulation results can be attributed to the same factors as in the case of self-guiding in a uniform-density plasma. Furthermore, the increase in electron energy and acceleration distance may be related to the mechanism described in Ref. 60: at later stages of the interaction, there is a significant overlap between the accelerated electrons and the red-shifted field of the laser pulse [see Fig. 5(a)], which may lead to the onset of direct laser acceleration [61, 62, 63]. This additional mechanism could enhance the final energy gain of the accelerated electrons, complementing the primary process of LWFA.
@@ VII Discussion on scaling
As shown in Sec. II, LWFA operated in the matched regime can be described using only two parameters, $n_{e}$ and $\tau_{0}$; alternatively, these parameters can be expressed in dimensionless form as $\mathcal{P}^{*}\coloneq\mathcal{P}_{0}/\mathcal{P}_{\mathrm{cr}}$ and $\tau^{*}\coloneq\tau_{0}\omega_{p}$. In general, $\mathcal{P}^{*}$ and $\tau^{*}$ may be functions that depend weakly on the laser energy. Treating them as constant for now, and disregarding the effects mentioned in the previous paragraph, we may derive the scaling of optimal LWFA input parameters, i.e., plasma density, pulse duration (or power), amplitude, spot size, and channel radius, all expressed in terms of laser energy:
@@
[…]
@@
[…]
@@
[…]
@@
[…]
$\frac{R_{\mathrm{ch}}}{\lambda_{0}}$ $=\frac{1.14}{\pi}\left(\frac{2\sqrt{\pi\ln{2}}}{\tau^{*}}\right)^{1/3}\left(\frac{\mathcal{E}_{0}}{\overline{\mathcal{E}}}\right)^{1/3}$ (34)
[…]
@@
[…]
@@
[…]
We see that the upper limit of the efficiency of energy transfer from the laser to the accelerated electron beam depends on neither the laser energy nor the wavelength,
$\eta=\frac{\mathcal{N}_{e}\mathcal{E}_{e,\mathrm{max}}}{\mathcal{E}_{0}}\approx 0.34\,\left(\tau^{*}\right)^{-1}\left(\mathcal{P}^{*}\right)^{1/6}.$ (38)
In the case of laser self-guiding in a uniform-density plasma (Sec. VI.1), the optimal case found in this work corresponds to $\mathcal{P}^{*}\approx 2.31$ and $\tau^{*}\approx 2.28$. Using Eqs. (29), (30), (32), and (33) to rescale the optimal parameters to, e.g., a $1\,\mathrm{kJ}$ laser pulse, we obtain $n_{e}\approx 1.56\times 10^{16}\,\mathrm{cm^{-3}}$, $\tau_{0}\approx 327\,\mathrm{fs}$, $a_{0}\approx 2.65$, and $w_{0}\approx 140\,\upmu\mathrm{m}$. For such a hypothetical LWFA setup, Eqs. (35)–(38) predict a maximum electron energy of $\approx 134\,\mathrm{GeV}$, an acceleration length of $\approx 9.8\,\mathrm{m}$, an electron charge of $\approx 1.2\,\mathrm{nC}$, and an acceleration efficiency of $17\,\%$, respectively.
In the case of laser guiding in a preformed plasma channel with matched radius (Sec. VI.2), the optimal case corresponds to $\mathcal{P}^{*}\approx 2.06$ and $\tau^{*}\approx 1.88$. Considering again a $1\,\mathrm{kJ}$ laser pulse, the optimal parameters translate to $n_{e}\approx 1.27\times 10^{16}\,\mathrm{cm^{-3}}$, $\tau_{0}\approx 299\,\mathrm{fs}$, $a_{0}\approx 2.54$, $w_{0}\approx 152\,\upmu\mathrm{m}$, and $R_{\mathrm{ch}}\approx 137\,\upmu\mathrm{m}$. Such a setup would yield an electron beam with a cut-off energy of $\approx 161\,\mathrm{GeV}$ over an acceleration distance of $\approx 11\,\mathrm{m}$, a charge of $\approx 1.3\,\mathrm{nC}$, and an efficiency of $20\,\%$.
Panels (a)–(f) of Fig. 6 show the scaling of optimal plasma density, laser pulse duration, laser power, beam waist, electron beam cut-off energy, and acceleration length, respectively, all expressed as functions of laser energy, according to Eqs. (29), (30), (31), (33), (35), and (36). Results are shown for both cases: laser self-guiding in a uniform-density plasma and guiding in a preformed plasma channel with matched radius.
@@ VIII Conclusion
To interpret the simulation results quantitatively, we derive novel analytical expressions for predicting the maximum electron energy and the corresponding acceleration length, accounting for the effects of laser pulse diffraction and energy depletion. The model shows relatively good agreement with the simulation results.
@@ Acknowledgements
Data availability
The data that support the findings of this article are openly available [75].
@@ Appendix A Simulation parameters
No. $n_{e}\left(10^{19}\mathrm{cm^{-3}}\right)$ $\tau_{0}\left(\mathrm{fs}\right)$ $w_{0}\left(\mathrm{\upmu m}\right)$ $a_{0}$ $\mathcal{E}_{e,\mathrm{max}}\left(\mathrm{MeV}\right)$ $l_{\mathrm{acc}}\left(\mathrm{\upmu m}\right)$
@@
Table 1: Parameters of PIC simulations with a uniform-density plasma. Electron density, $n_{e}$, pulse duration, $\tau_{0}$, beam waist, $w_{0}$, and strength parameter, $a_{0}$, prescribed in PIC simulations $A_{1}$ - $A_{18}$ as well as the corresponding cut-off electron energy $\mathcal{E}_{e,\mathrm{max}}$ acquired over acceleration distance $l_{\mathrm{acc}}$.
No. $n_{e}\left(10^{19}\mathrm{cm^{-3}}\right)$ $\tau_{0}\left(\mathrm{fs}\right)$ $w_{0}\left(\mathrm{\upmu m}\right)$ $a_{0}$ $R_{\mathrm{ch}}\left(\mathrm{\upmu m}\right)$ $\mathcal{E}_{e,\mathrm{max}}\left(\mathrm{MeV}\right)$ $l_{\mathrm{acc}}\left(\mathrm{\upmu m}\right)$
@@
Table 2: Parameters of PIC simulations with a preformed plasma channel of matched radius. Electron density, $n_{e}$, pulse duration, $\tau_{0}$, beam waist, $w_{0}$, strength parameter, $a_{0}$, and the channel radius, $R_{\mathrm{ch}}$, prescribed in PIC simulations $B_{1}$ - $B_{18}$ as well as the corresponding cut-off electron energy $\mathcal{E}_{e,\mathrm{max}}$ acquired over acceleration distance $l_{\mathrm{acc}}$.